OpenStudy (anonymous):

Find a polynomial of degree 4 with -3 as a zero of multiplicity 2 and 0 and 6 as zeros of multiplicity 1

OpenStudy (zzr0ck3r):

so we have the things we need (x+3) will give us a zero at -3 (x-0) = x will give us a zero at 0 (x-6) will give us a zero at 6 we want the -3 one to be of multiplicity 2 so (x+3)^2 we want the other two to have multiplicity 1 x^1 and (x+6)^1 so our polynomial is \[\begin{equation} \begin{aligned} x(x+3)^2(x-6) &=x[(x^2+6x+9)(x-6)] \\ &=x(x^3-6x^2+6x^2-36x+9x-54) \\ &=x(x^3-29x-54) \\ &=\large \color{red}{x^4-29x^3-54x} \end{aligned} \end{equation} \]

OpenStudy (zzr0ck3r):

easy peasy

OpenStudy (anonymous):

thanks @zzr0ck3r